(6x)4 when x=2
the four is to the fourth power. i do not want the answer but i need to now if i put six and x to the fourth power then multiply or if i multipy 6 and x then fourth power.

Answer by ankor@dixie-net.com(15659)

(6x)^4 when x=2
the four is to the fourth power.
:
Both 6 and x have to raised to the 4th power
:
You can replace x with 2
(6*2)^4 = 12^4
:
or you can write it
6^4 * x^4 = 1296x^4
:
Replace x with 2
1296*2^4 = 1296*16
: